OpenStudy (anonymous):

Need help if you give me the answer , I'll give you a medal :D --> Solve 3x-11=-41

OpenStudy (anonymous):

3x - 11 = -41 add 11 to both sides 3x - 11 + 11 = -41 + 11 3x = -30 divide both sides by 3 x = -10